OpenStudy (anonymous):

The height h (in feet) of an object shot into the air from a tall building is given by the function h(t)=800+80t-16t^2, where t is the time elapsed in seconds. What should be the velocity when it reaches its highest point? when does this occur? And what is the maximum height reached by the object?

OpenStudy (anonymous):

i know the velocity is v(t)=80-32t

OpenStudy (anonymous):

I just don't know where to go from there.

OpenStudy (anonymous):

velocity should be zero as the object is stationary for an instant before it begins to fall

OpenStudy (anonymous):

the highest point is at the max value of the height functn

OpenStudy (anonymous):

ur derivative of the height function appears correct

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

so different s(t) wrt t, and find the value of time from there by equating it to zero

OpenStudy (anonymous):

so I would just set the velocity equal to zero, solve for x then plug it into the position function?

OpenStudy (anonymous):

yes..that gives you the maximum height

OpenStudy (anonymous):

do you think its appropriate to answer the second question as "right before the ball begins to fall"?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok wonderful! thank you!

OpenStudy (anonymous):

Set \(80-32t=0 \implies t=\frac{5}{2}\). So, at \(t=\frac{5}{2}\), the object reaches its maximum hieght. Its velocity at that point must be \(0\) and its hieght is \(800+80\frac{5}{2}-16(\frac{5}{2})^2\).

OpenStudy (anonymous):

thanks guys.

OpenStudy (anonymous):

i have to put my two cents in sorry. you do not need any calculus for this. the vertex of a quadratic polynomial is given by \[-\frac{b}{2a}\] and in this case you get \[\frac{80}{32}=\frac{5}{2}\] just thought i would mention it now i will be quiet

OpenStudy (anonymous):

thanks guys.